# This work has been submitted to the IEEE Access for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible. xx.xxxx/ACCESS.2024.xxxxxxx

### v2

@@ III-A Overview of SNNs
Spiking Neural Networks (SNNs) mimic the brains’ functionality through the utilization of spikes for transmitting information [37, 266, 267]. Therefore, SNNs encode information into discrete spike trains. Popular encoding techniques include rate coding [268] (e.g., spike count, spike density, population activity [269]), and temporal coding (e.g., burst [270], time-to-first spike (TTFS) [271], phase [272], and rank-order [273]). Each spiking neuron processes the input spikes, and its behavior (i.e., neuronal dynamics) depends on the neuron model, such as Hodgkin–Huxley (HH) [274], Leaky Integrate and Fire (LIF) [275], Izhikevich Model [276], and Adaptive Exponential Integrate-and-Fire (AdExIF) [277]. Neuron model selection typically considers the expected neuronal dynamics and computational complexity [278, 279]. The output spikes are generated only when neurons’ membrane potential reaches the threshold, and transmitted through synapses, enabling ultra-low power/energy consumption [52, 280, 281, 282, 283].
@@
SNN training can be performed using bio-plausible or analytical learning rules. The bio-plausible ones are typically characterized by local learning mechanism, like Hebbian [284], Spike-Timing-Dependent Plasticity (STDP) [285], Spike-Driven Synaptic Plasticity (SDSP) [286, 287], and Reward-Modulated STDP (R-STDP) [288, 289]. Meanwhile, the analytical ones encompass DNN-to-SNN conversion [290], Surrogate Gradient Learning [291] (e.g., Backpropagation Through Time (BPTT) [291, 292, 293], Spatio-Temporal Backpropagation (STBP) [294], and Deep Continuous Local Learning (DECOLLE) [295]), Spike-Triggered Local Representation Alignment (ST-LRA)  [296], and Bayesian Learning [297].
@@
SNN processing demands a suitable computing hardware to maximize its potentials in accuracy, latency, and power/energy efficiency. Conventional von-Neumann architecture-based hardware platforms (e.g., CPU and GPUs) have been widely used to perform Python-based SNN processing using generic arithmetic units  [298, 299, 300, 301], thereby leading to sub-optimal efficiency gains. To address this, CMOS-based neuromorphic hardware accelerators facilitate efficient spike transmission and computation [302], allowing their implementation in the Field-Programmable Gate Array (FPGA) or Application-Specific Integrated Circuit (ASIC). Popular accelerators include Neurogrid [303], ROLLS [304], TrueNorth [305], and Loihi [306]. Beyond CMOS-based technologies, the processing-in-memory (PIM) or compute-in-memory (CIM) paradigm has been explored to further reduce latency and energy in data transfer between memory and compute units [307, 308] by leveraging non-volatile memory (NVM) technologies, such as Resistive Random Access Memory (RRAM), Magnetic RAM (MRAM), and Phase Change Memory (PCM) [309].
III-B Neuromorphic Continual Learning (NCL)
NCL is an advancement of bio-plausible SNNs for facilitating continual adaptation in dynamic environments. Here, we discuss the state-of-the-art NCL methods, including their key ideas and limitations, whose details are summarized in Table IV.
III-B1 Enhancements on Unsupervised STDP Learning
@@
While conventional STDP enables efficient online unsupervised learning, employing STDP alone may still suffer from CF [41, 42]. To address these challenges, proposed techniques have been proposed in the literature, as discussed in the following.
@@
Adaptive Synaptic Plasticity (ASP) [97]: It integrates weight decay with the STDP-based weight updates to balance forgetting and learning, while leveraging time-dependent learning rate. ASP uses a two-phase weight update process (i.e., recovery and decay), allowing weight updates based on spiking activities and gradually leak toward a baseline value; see Fig. 13. Despite its strengths, ASP requires larger quantities of input samples from earlier distributions than later ones, hence requiring the knowledge of task changes [96], making it unsuitable for OCL scenario.
@@
Controlled Forgetting Network (CFN) [96]: It is an SNN architecture that exploits dopaminergic neurons to modulate the synaptic plasticity. Its idea is to temporarily make the weights of some neurons more plastic (easier to change) and keep the weights of other neurons. The modulation is triggered by dopaminergic neurons, which fire when there are no or only a small number of incoming inhibitory spikes (indicating that new information is encountered), and then stimulate a temporary boost of learning rate for other neurons (Fig. 14). Despite its benefits, CFN requires additional components (i.e., dopaminergic neurons and their connections) and still considers a simple MNIST dataset with conventional CL scenario, hence making it challenging for deployments with more complex datasets under OCL scenario.
@@
SpikeDyn [41]: It focuses on enhancing STDP for enabling unsupervised CL while minimizing energy consumption in both training and inference phases (Fig. 15). Its key ideas include: reduction of neuronal operations by substituting the inhibitory neurons with direct lateral inhibitions to reduce memory and energy requirements; enhancing unsupervised CL algorithm by employing adaptive learning rates, weight decay, adaptive threshold potential, and reduction of spurious weight updates; and SNN model size search by leveraging analytical models to estimate the memory and energy requirements, and selecting a Pareto-optimal model that meets the resource constraints. SpikeDyn demonstrated that optimization techniques can be exploited to minimize memory and energy requirements of CL systems, but they should be supported with an enhanced unsupervised CL algorithm to maintain the performance.
@@
lpSpikeCon [42]: It extends the studies in SpikeDyn [41], by enabling STDP-based unsupervised CL under low-precision settings for autonomous embedded systems (e.g., robots). lpSpikeCon employs weight quantization and compensates the loss of information by identifying and adjusting SNN parameters that significantly impact accuracy (Fig. 16). This study found that the key SNN parameters to adjust include the adaptive threshold potential and weight decay rate. This study demonstrated that low-precision settings can be exploited for substantially reducing memory and energy requirements, and adjustments of other parameters are crucial to maintain the performance.
@@
Despite their benefits, SpikeDyn and lpSpikeCon still consider a simple MNIST dataset and conventional CL scenario, thus requiring further studies for OCL scenario.
III-B2 Predictive Coding

### v3

@@ III-A Overview of SNNs
A neuromorphic computing system encompasses SNN processing (including spike-based operations and training) in the software part, and a neuromorphic processor in the hardware part. The overview of the system is illustrated in Fig. 13, and described in the following, answering key question Q4.
Figure 13: An overview of a neuromorphic system, encompassing hardware (HW) and software (SW) parts.
@@
SNN Models and Operations: SNNs mimic the brains’ functionality through the utilization of spikes for transmitting information [38, 266, 267]. Therefore, SNNs encode information into discrete spike trains. Popular encoding techniques include rate coding [268] (e.g., spike count, spike density, population activity [269]), and temporal coding (e.g., burst [270], time-to-first spike (TTFS) [271], phase [272], and rank-order [273]). Each spiking neuron processes the input spikes, and its internal behavior (i.e., neuronal dynamics) depends on the spiking neuron model, such as Hodgkin–Huxley (HH) [274], Leaky Integrate and Fire (LIF) [275], Izhikevich Model [276], and Adaptive Exponential Integrate-and-Fire (AdExIF) [277]. Neuron model selection typically considers the expected neuronal dynamics and computational complexity [278, 279]. The illustration of neuronal dynamics and computational complexity trade-off for different neuron models is shown in Fig. 14. The output spikes are generated only when neurons’ membrane potential reaches the threshold, and transmitted through synapses, enabling ultra-low power/energy consumption [50, 280, 281, 282, 283].
Note, spike-based operations can also be applied to emerging network models like Transformers, i.e., so-called Spike-based Transformers, by employing spike data representation and spiking neuron model in Transformer network architectures, as demonstrated in recent research works [284, 285, 286, 287, 288]. These works show that Spike-based Transformers have comparable scalability and performance over the conventional Transformers, while enabling higher energy efficiency when executed on neuromorphic hardware due to their sparse spike-based operations.
Figure 14: Trade-off between bio-plausibility and computation complexity for different neuron models; adapted from studies in [276].
@@
SNN Training: The training process can be performed using bio-plausible or analytical learning rules. The bio-plausible ones are typically characterized by local learning mechanism, like Hebbian [289], Spike-Timing-Dependent Plasticity (STDP) [290], Spike-Driven Synaptic Plasticity (SDSP) [291, 292], and Reward-Modulated STDP (R-STDP) [293, 294]. Meanwhile, the analytical ones encompass DNN-to-SNN conversion [295], Surrogate Gradient Learning [296, 297, 298] (e.g., Backpropagation Through Time (BPTT) [296, 299, 300], Spatio-Temporal Backpropagation (STBP) [301], and Deep Continuous Local Learning (DECOLLE) [302]), Spike-Triggered Local Representation Alignment (ST-LRA)  [303], and Bayesian Learning [304]. To facilitate SNN training, software frameworks play a vital role. Some notable ones include SpikingJelly [305], BindsNet [306], SNNtorch [307], TinySpiking [308], Norse [309], SpyTorch [310], SINABS [311], SpykeTorch [312], Brian [313], Lava [314], BrainCog [315], PySNN [316], and PyNCS [317]. These frameworks are still evolving to support CL features. Meanwhile, standard DL frameworks such as Tensorflow [318] and PyTorch [319], and CL frameworks such as PyCIL [320], FACIL [47] and Avalanche [321] stand out for their user-friendly interfaces, flexibility, and support for complex applications. Table VII addresses the key question Q5 by comparing neuromorphic, DL and CL frameworks, outlining that the SNN frameworks with local STDP learning and neuromorphic hardware are more computationally efficient.
TABLE VII: Comparison of neuromorphic, standard DL and CL frameworks.
Frameworks Learning Mechanism CL Support Hardware Support Flexibility and Usability PyTorch-based Computational Efficiency
Unsupervised Supervised Hybrid
Neuromorphic Frameworks
SpikingJelly [305] STDP, Hebbian Learning BPTT, SG RL CPU, GPU, Neuromorphic chips High High
BindsNet [306] STDP, Hebbian Learning $\times$ RL $\times$ CPU, GPU Medium Moderate
SNNtorch [307] STDP BPTT, SG $\times$ $\times$ CPU, GPU High High
TinySpiking [308] STDP $\times$ $\times$ $\times$ CPU, MCU High Python-Based High
Norse [309] STDP, Hebbian Learning BPTT, SG RL $\times$ CPU, GPU High High
SpyTorch [310] STDP $\times$ $\times$ $\times$ CPU, GPU Medium Moderate
SINABS [311] STDP BP $\times$ $\times$ Neuromorphic chips Medium High
SpykeTorch [312] STDP, R-STDP $\times$ $\times$ $\times$ CPU, GPU Medium Python-Based Moderate
Brian[313] STDP, Hebbian Learning $\times$ $\times$ $\times$ CPU Medium Python-Based Moderate
Lava [314] STDP, Hebbian Learning $\times$ RL $\times$ Neuromorphic chips Medium Python-Based High
BrainCog [315] STDP, Hebbian Learning BPTT, SG RL CPU, GPU, Neuromorphic chips High High
PySNN[316] STDP SG $\times$ $\times$ CPU, GPU Medium High
PyNCS [317] STDP, Hebbian Learning $\times$ $\times$ $\times$ Neuromorphic chips Medium Python-Based Moderate
Standard DL Frameworks
TensorFlow [318] $\times$ BP RL CPU, GPU, TPU Very High Python-Based Relatively low
PyTorch [319] $\times$ BP RL CPU, GPU Very High Relatively low
CL Frameworks
PyCIL [320] $\times$ BP RL CPU, GPU High Relatively low
FACIL[47] $\times$ BP, SG $\times$ CPU, GPU High Relatively low
Avalanche [321] $\times$ BP RL CPU, GPU High Relatively low
@@
Neuromorphic Hardware: SNN processing demands a suitable computing hardware to maximize its potentials in accuracy, latency, and power/energy efficiency. Conventional von-Neumann architecture-based hardware platforms (e.g., CPU and GPUs) have been widely used to perform Python-based SNN processing using generic arithmetic units  [322, 306, 305, 323], thereby leading to sub-optimal efficiency gains. To address this, CMOS-based neuromorphic hardware accelerators facilitate efficient spike transmission and computation [324], allowing their implementation in the Field-Programmable Gate Array (FPGA) or Application-Specific Integrated Circuit (ASIC). Popular accelerators include Neurogrid [325], ROLLS [326], TrueNorth [327], Loihi [328], DYNAP [329], and Akida [330][331]. Their typical hardware architecture is shown in Fig. 15(a). Beyond CMOS-based technologies, the processing-in-memory (PIM) or compute-in-memory (CIM) paradigm has been explored to further reduce latency and energy in data transfer between memory and compute units [332, 333] by leveraging non-volatile memory (NVM) technologies, such as Resistive Random Access Memory (RRAM), Magnetic RAM (MRAM), and Phase Change Memory (PCM) [334]. Their typical hardware architecture is shown in Fig. 15(b).
Figure 15: Neuromorphic processors: (a) CMOS-based HW architecture, and (b) NVM-based HW architecture; adapted from studies in [323].
III-B Framework, Datasets, Benchmarks, and Evaluation Metrics
The discussion in this section addresses the key questions Q4 and Q8.
Framework: Recently, the SNN community has been building a framework for benchmarking neuromorphic algorithms and systems, so-called NeuroBench [335], encompassing datasets and metrics. Its algorithm-level metrics include correctness (e.g., mean average precision (mAP)) and complexity (e.g., memory footprint, synaptic operations, as well as model and activation sparsity). Meanwhile, its system-level metrics include task-specific correctness, timing, and efficiency. Despite these advancements, NeuroBench framework is still under continuous development, and has not provided comprehensive benchmarks for NCL considering different CL settings and scenarios (see Section II-C and Section II-E). Therefore, a standardized framework for comprehensively benchmarking NCL is still missing, and requires extensive developments. Toward this, we propose to consider the following additional datasets, benchmarks, and evaluation metrics.
Datasets & Benchmarks: Neuromorphic datasets play a crucial role in advancing the NC field, particularly in the training and evaluation of SNN models. These datasets are derived or generated to leverage the unique characteristics of event-based data. Some datasets are obtained using neuromorphic sensors, and some others are derived from existing conventional datasets. Neuromorphic datasets can be divided into three categories based on their acquisition methods, as the following.
  1. 1.
    Real-World Scene Datasets: They are collected directly from real-world environments using event-based sensors, such as Dynamic Vision Sensor (DVS) cameras. Consequently, these datasets are unlabeled until the labeling process is performed. For instance, the DVS-128 Gesture dataset [336] for gesture recognition is obtained using a DVS camera.
  2. 2.
    Transformation Datasets: They are obtained by transforming the existing (labeled) datasets from the ANN domain through recordings with event-based sensors. An example for this category is the CIFAR10-DVS dataset [337]. These datasets are expected to leverage event-based data characteristics for SNNs, while having direct counterparts in the ANN domain for comparison, thereby making them popular for SNN evaluation.
  3. 3.
    Algorithmically Generated Datasets: They are synthesized using algorithms that emulate the behavior of event-based sensors. They convert existing labeled image, video, or speech data into neuromorphic datasets using difference-based algorithms. An example for this category is the Spike-TIDIGITS [338].
Table VIII provides an elaborate comparison of the prominent neuromorphic datasets, including their publication year, characteristics, and tasks. Currently, they also serve as benchmarks for evaluating and comparing SNN models, including the NCL methods. However, these datasets are mainly for classification tasks, which limit the applicability of neuromorphic systems. Therefore, more diverse datasets are required including the types of tasks (e.g., regression and generative), diversity in conditions during data collection (e.g., weather, lighting, fog, etc.), and even diversity of application use-cases (e.g., vision, sound, tactile, and olfactory datasets). Moreover, benchmarks for NCL should also consider different possible CL settings (e.g., supervised, unsupervised, and RL) and scenarios (e.g., task-IL, domain-IL, and class-IL).
Evaluation Metrics: Section II-D highlights the standard evaluation metrics including average accuracy, BWT and FWT, forgetting measure, model size, sample size, compute resources (e.g., FLOPS and time complexity), and power/energy efficiency. Therefore, they cover limited metrics for evaluating the functionality of different CL tasks. Moreover, the conventional metrics serve as indicators of a model’s ability to retain prior knowledge while learning new tasks sequentially. However, when applied to neuromorphic systems, these metrics must be reinterpreted and complemented due to the distinct computational paradigms inherent in such architectures, such as event-driven computation, quantized activations, low-precision operations, and temporal spike-based encoding. To address this, we reference works that propose complementary performance indicators and suggest new metrics as follows.
  • Performance: In spike-based systems, processing time often relies on the neural coding (e.g., rate, TTFS, etc.). Therefore, latency (measured in time steps until a decision is made) and throughput becomes a critical indicator of performance of low-power or real-time applications [339].
  • Efficiency: Power and energy consumption during CL training and inference are important to indicate the energy efficiency of NCL. For estimation, spike count and membrane potential dynamics may also provide crucial insights into the energy efficiency [42]. Furthermore, utilization of memory bandwidth and compute module, as well as synaptic operation rates (SOPS) are also relevant for indicating efficiency when running networks on neuromorphic chips. NCL systems that maintain accuracy with high efficiency are preferred for embedded AI deployments.
  • Robustness: Neuromorphic systems may exhibit variability due to low-bit precision, noisy spikes, and device-level process variations. Therefore, robustness (defined as accuracy maintained under noise, perturbations, or quantization) emerges as a meaningful metric besides accuracy.
  • Task-specific functionality metrics: They depend on the type of task. For instance, mean absolute error (MAE) and mean square error (MSE) can be used for regression-based tasks.
  • Memory ratio: It represents how much memory overhead compared to the original model size, that is needed to execute the implemented CL method.
  • Adaptability scores: They represent how effective is the implemented CL method over time considering different CL scenarios i.e., task-IL, class-IL, domain-IL, TA etc.
TABLE VIII: Comparison of widely used neuromorphic datasets for SNNs with their publication year, description, key features and tasks.
Dataset Year Description Key Features Tasks
N-MNIST 2015 An event-based version of the MNIST Classes: 10 Digit Classification
[340] dataset captured using Asynchronous Resolution: 28x28
Time-based Image Sensor (ATIS) Training Samples: 60,000
with each digit recorded over 300 ms Testing Samples: 10,000
DVS-128 Gesture 2017 An event-driven data of hand and arm Classes: 11 Gesture Recognition
[336] gestures captured using DVS128 camers, Subjects: 29
focusing on dynamic movement with Illumination conditions: 3
6s/sample duration Resolution: 128x128
CIFAR10-DVS 2017 An event-based version of CIFAR-10 Classes: 10 & Event streams: 10,000 Image Classification
[337] dataset with 300ms/sample duration Resolution: 128x128
N-Caltech101 2015 A spiking version of the original Samples: 8709 Object Recognition
[340] frame-based Caltech101 dataset Classes: 100 objects, and
with 300–500ms/sample duration a background class
N-CARS 2018 A large real-world event-based dataset Classes: 2 Car & Background
[341] captured using ATIS camera with Samples of Car: 12,336 & Background: 11,693 Recognition
100ms/sample duration Training Samples: 7940 (car), 7482 (bg)
Testing Samples: 4396 (car), 4211 (bg)
N-Omniglot 2022 A large-scale neuromorphic dataset Classes: 1,623 Few-Shot Learning
[342] of handwritten characters, representing Samples per class: 20 in SNNs
50 different languages captured by DVS Samples: 32,460
Spiking Heidelberg 2022 An audio-based dataset of 10k high- Classes: 10 Speech Classification
Digits (SHD) [343] quality recordings of spoken digits ranging Digit Count: 10,420 and Keyword Spotting
from zero to nine in English and German Speakers: 12
ES-ImageNet 2021 A large-scale event-stream dataset Classes: 1000 Image Classification
[344] converted from the ImageNet dataset [345] Training Samples: 1,257,035
using a software-based event generation Testing Samples: 49,881
Omnidirectional Discrete Gradient (ODG) Total Samples: 1.3 M
algorithm with 29.47ms/sample duration Resolution: 224×224*
SpikeBALL 2023 A neuromorphic dataset capturing 10 Class: 1 Object Tracking
[346] different trajectories of a ball in a table Events: 80,000 to 100,000
football game per trajectory
NE15-MNIST 2016 A spiking version of MNIST dataset Classes: 10 Visual Recognition
[347] with four sub-datasets encoded using: Training Samples: 60,000
Poissonian, Rank-Order, DVS recorded Testing Samples: 10,000
flashing and DVS recorded moving
ASL-DVS 2019 Event-driven data of 24 letters (A-Y, excl. J) Classes: 24 Sign Language
[348] from American sign language recorded Samples: 100,800 / Samples per-class: 4,200 Recognition
using DAVIS240c NVS camera Training Samples: 80,640
with 100ms/sample duration Testing Samples: 20,160
N-TIDIGITS18 2018 The spike recordings from a Dynamic Classes: 11 Spoken Digit
[349] Audio Sensor (DAS) in response to the Training Samples: 8,623 Recognition [350]
TIDIGITS audio dataset Testing Samples: 8,700
Spike-TIDIGITS 2020 A Spike-based version of TIDIGITS dataset Classes: 11 Speech Recognition
[338] created using the Biologically Plausible Training Samples: 2464
Auditory Encoding (BAE) algorithm Testing Samples: 2486
Spike-TIMIT [338] 2020 An event-based version of TIMIT Training Samples: 4621 Audio Classification
dataset [351] Testing Samples: 1679
DVS Lane Extraction 2019 A high-resolution event-based dataset Classes: 5, (0 for bg and Lane Extraction
(DET) [352] of complex traffic scenes and various lane 1,2,3,4 for four lane types)
types annotated with multi-class Images: 5,424
segmentation captured by CeleX-V DVS Resolution: 1280×800
KUL-UAVSAFE [353] 2021 A joint DVS and RGB dataset in an indoor Subjects: 6 & Variations: Clothes color, People Detection
environment with one walking human shape & walking style
  • *: Events are generated within a 256x256-pixel range, but only those in the central 224x224-pixel region are meaningful, the rest are noise caused by edge motion; bg: background.
III-C Neuromorphic Continual Learning (NCL)
The integration of CL with neuromorphic computing (i.e., NCL) is the central focus of the paper. Specifically, how NCL facilitates continual adaptation in diverse operational environments by leveraging key characteristics of neuromorphic systems that align with CL requirements, as highlighted in the following.
  • Neural plasticity: It is established using bio-plausible learning rules, such as Hebbian rule and Spike-Timing Dependent Plasticity (STDP), which enable SNNs to flexibly learn new information while preserving prior knowledge (e.g., through weight potentiation and depression), which is useful for addressing CF problem.
  • Localized learning in each synapse: Bio-plausible learning rules in SNNs can leverage spiking activities for updating weight locally in each synapse. It is useful for learning information without explicit labels, which is required for enabling online training in updating systems’ knowledge and adapting to dynamic environments.
  • Event-based processing: The event-driven operations in SNNs consider both spatial and temporal information in the event-based data streams, which aligns with the nature of dynamic data streams in CL. This potentially streamlines the execution of CL algorithms.
  • Energy-efficient computation: The event-driven operations in SNNs and neuromorphic hardware enable ultra-low power/energy computation, which is required for enabling CL execution in tightly-constrained systems.
Figure 16: A state-of-the-art taxonomy of NCL methods covered in this survey. We have highlighted the main categories (blue blocks), with each of their works shown (red blocks).
III-D NCL Methods
In this sub-section, we provide an in-depth review of state-of-the-art NCL methods, including their key ideas and techniques, whose details are summarized in Table IX. Fig. 16 presents the taxonomy that systematically categorizes these methods based on their specific techniques. This discussion also addresses the key question Q6.
III-D1 Enhancements on Unsupervised STDP Learning
@@
While conventional STDP enables efficient online unsupervised learning, employing STDP alone may still suffer from CF [42, 43]. To address these challenges, several techniques have been proposed in the literature, as discussed in the following.
@@
Adaptive Synaptic Plasticity (ASP) [102]: It integrates weight decay with the STDP-based weight updates to balance forgetting and learning, while leveraging time-dependent learning rate. ASP uses a two-phase weight update process (i.e., recovery and decay), allowing weight updates based on spiking activities and gradually leak toward a baseline value; see Fig. 17. The SNN was implemented in Brian [313], (a lightweight SNN simulator for rapid prototyping and experimentation with a flexible Python interface), to perform dynamic digit recognition with the MNIST dataset. The digit categories ’0’ to ’9’ were presented sequentially, without intermixing at any point during the training phase (i.e., no data reinforcement). For a 6400 neuron SNN, ASP achieved avg. accuracy of 94.2% outperforming the standard STDP. A summary is presented in Table IX and the empirical results are shown in Table XI and Table XII, providing a comparative analysis in terms of network architecture, size and key performance metrics such as accuracy, memory footprint, and power/energy consumption of the reviewed works. Despite its strengths, ASP requires larger quantities of input samples from earlier distributions than later ones, hence requiring the knowledge of task changes [101], making it unsuitable for OCL scenario.
@@
Controlled Forgetting Network (CFN) [101]: It is an SNN architecture that exploits dopaminergic neurons to modulate the synaptic plasticity. Its idea is to temporarily make the weights of some neurons more plastic (easier to change) and keep the weights of other neurons. The modulation is triggered by dopaminergic neurons, which fire when there are no or only a small number of incoming inhibitory spikes (indicating that new information is encountered), and then stimulate a temporary boost of learning rate for other neurons (Fig. 18). For a 6,400 neuron CFN, the evaluations on MNIST dataset in a fully disjoint scenario, achieved on average 95.24% classification accuracy across all digits (see summary in Table IX and the comparative empirical analysis in Table XI and Table XII). Despite its benefits, CFN requires additional components (i.e., dopaminergic neurons and their connections) and considers a conventional CL scenario (i.e., temporally separated tasks), which makes it challenging for deployments in an environment of changing priorities, i.e., OCL scenario.
@@
SpikeDyn [42]: It focuses on enhancing STDP for enabling unsupervised CL while minimizing energy consumption in both training and inference phases (Fig. 19). Its key ideas include: reduction of neuronal operations by substituting the inhibitory neurons with direct lateral inhibitions to reduce memory and energy requirements; enhancing unsupervised CL algorithm by employing adaptive learning rates, weight decay, adaptive threshold potential, and reduction of spurious weight updates; and SNN model size search by leveraging analytical models to estimate the memory and energy requirements, and selecting a Pareto-optimal model that meets the resource constraints. The SNN was implemented using Bindsnet, a ML-oriented SNN library in Python [306]. Evaluations on MNIST dataset in a dynamic environment (where tasks are fed sequentially, without re-feeding earlier ones, and each task contains an equal number of samples), for a 200 neuron SNN, showed avg. 23% and 4% improved accuracy than ASP [102] when learning a new task and when retaining the old task, respectively. It also reduced energy consumption by up to avg. 57% and 51% for training and inference, respectively (see summary in Table IX and the comparative empirical analysis in Table XI and Table XII). SpikeDyn demonstrated that optimization techniques can be exploited to minimize memory and energy requirements of CL systems, but they should be supported with an enhanced unsupervised CL algorithm to maintain the performance.
@@
lpSpikeCon [43]: It extends the studies in SpikeDyn [42], by enabling STDP-based unsupervised CL under low-precision settings for embedded AI systems (e.g., robots). lpSpikeCon employs weight quantization and compensates the loss of information by identifying and adjusting SNN parameters that significantly impact accuracy (Fig. 20). The SNN was implemented using Bindsnet [306], see Table VII. It performed a case study in a dynamic CL scenario using the MNIST dataset. The network was trained sequentially on digit classes from 0 to 9, receiving one class at a time. After each training phase, the model was evaluated on the test samples of all classes learned so far, simulating a real-time CL process. The 4-bit weight quantized SNN with key parameter adjustments showed no accuracy loss in the inference while reducing the weight memory by 8x compared to 32-bit non-quantized SNN (see summary in Table IX and comparative empirical analysis in Table XI and Table XII). This study found that the key SNN parameters to adjust include the adaptive threshold potential and weight decay rate. This study demonstrated that low-precision settings can be exploited for substantially reducing memory and energy requirements, and adjustments of other parameters are crucial to maintain the performance. Despite their promising results, SpikeDyn and lpSpikeCon have so far focused on the MNIST dataset and conventional CL scenario. Expanding these methods to more complex scenarios for OCL is a potential future research direction.
@@
III-D2 Predictive Coding